OpenStudy (anonymous):

Short Answer: Raise the quantity in parentheses to the indicated exponent, and simplify the resulting expression. Express answers with positive exponents. (28xy^3/130x^3y^-2)^3

OpenStudy (anonymous):

Jimmy do you think you can help me

OpenStudy (anonymous):

ill be back - just been called away

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

sorry everyone the problem is wrong let me repost it k

OpenStudy (cathyangs):

oh...this is messy! When you raise an exponent to an exponent, like that 3y^-2, you have to treat it like 'normal' powers. that power would simplify to 1/9x^2 which is the power that you're raising 130x to.

OpenStudy (anonymous):

(28xy^3/140x^3y^-2)^2 is the actual problem

OpenStudy (cathyangs):

ok...my previous reply still stands.

OpenStudy (anonymous):

don't you have to multiply the exponents to get the right answer or do you add them?

OpenStudy (anonymous):

wouldn't it look like thing 28*28 x*x y^3*y^3/140*140*x^3*x^3y^-2y^-2 is how the problem really looks right

OpenStudy (cathyangs):

no, you multiply them like here:\[(x ^{3})^{4}\] this would where it simplifies to \[x ^{?}\] you add them here:\[x ^{2}*x ^{3}\] which equals \[x ^{5}\]

OpenStudy (cathyangs):

not quite, you have to simplify the exponents in the denominator, get it down to one "layer"

OpenStudy (anonymous):

how do I do that

OpenStudy (cathyangs):

I said before, up there^

jhonyy9 (jhonyy9):

Sinz do you know that y on exponent -2 = 1/y squared yes ?

OpenStudy (anonymous):

so basically break the equation down to this 28/140 x/x^3 y^3/y^-2 all by ^2 right

OpenStudy (cathyangs):

I said before, up there^

OpenStudy (cathyangs):

x^-a is equal to 1/x^a

OpenStudy (cathyangs):

I said before, up there^

jhonyy9 (jhonyy9):

you need correcting this cathy

OpenStudy (anonymous):

okay

jhonyy9 (jhonyy9):

Sinz do you know the result now sure ?

OpenStudy (cathyangs):

oops...x^3/x^2=x^1

jhonyy9 (jhonyy9):

yes this was

OpenStudy (cathyangs):

was that what you were talking about?

jhonyy9 (jhonyy9):

yes because you have wrote there x squared divide x squared = x

OpenStudy (anonymous):

still a little confused but trying to figure it out also on paper to see how it works

jhonyy9 (jhonyy9):

ok write it here please

OpenStudy (anonymous):

so I have to use the PEMDAS law to solve this problem do everything in the parenthesis and go from there

OpenStudy (cathyangs):

yes..I guess.

jhonyy9 (jhonyy9):

after you will can understanding it how you need to solve it will see that is very easy

OpenStudy (anonymous):

(28xy^3/140x^3y^-2)^2 this is what I am trying to solve

OpenStudy (anonymous):

so if that's the case then I would sit there and double everything in the parenthesis first right

jhonyy9 (jhonyy9):

yes i want you say that first you need to make that calcule inside parantheses

OpenStudy (cathyangs):

so simplify the stuff inside parenthesis, and such. not doubling, taking it all to the power of two, multiplying all that by all that.

OpenStudy (anonymous):

okay so it is 28^2 x^2 y^5/140^2 x^5y^4 right?

OpenStudy (cathyangs):

the denominator should be 140^2 x^2/(9x^2, I believe

OpenStudy (anonymous):

my answer would look like this 784y/x^3 is that right

OpenStudy (anonymous):

sorry miss calculated it is .14y/x^3

jhonyy9 (jhonyy9):

not is correct

OpenStudy (cathyangs):

oops...I've been reading the problem the wrong way. I got y^10/25x^4

OpenStudy (cathyangs):

jhonnyy9, is english your second language?

OpenStudy (anonymous):

jhony how would you do the problem could you help

jhonyy9 (jhonyy9):

yes cathy but your last answer is the right answer sure

OpenStudy (cathyangs):

XD you think my name is cathy? (I know most people abbreviate screen names here)

OpenStudy (anonymous):

but how did she get the answer that is the question

jhonyy9 (jhonyy9):

sorry

jhonyy9 (jhonyy9):

Sinz

OpenStudy (anonymous):

??

jhonyy9 (jhonyy9):

first you need calcule inside parantheses

jhonyy9 (jhonyy9):

so 28/140 = 1/5

OpenStudy (anonymous):

or .2

jhonyy9 (jhonyy9):

x/x3 = 1/x2

OpenStudy (cathyangs):

I'm a she now!!! :D sorry...off topic I worked it all out on paper, I first simplified the denominator, the bottom half of the fraction. then I wrote it out and multiplied all the like terms.

OpenStudy (cathyangs):

yes, you have to simplify the parentheses completely, first.

jhonyy9 (jhonyy9):

and y on exponent -2 = 1/y2 but because this is in denominator so than y2 will being part of numinator

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

okay

jhonyy9 (jhonyy9):

and at finaly

OpenStudy (cathyangs):

you don't even have to write y^-2 as 1/y^2, because it goes away when you simplify the ys in the fraction.

jhonyy9 (jhonyy9):

you see now why will be the answer y10/25x4

jhonyy9 (jhonyy9):

cathy no you not can simplify the y from the fraction never

OpenStudy (anonymous):

okay so you had to add the y^-2 to the numerator right which made it y^5?

jhonyy9 (jhonyy9):

so because there is y3/y on -2 this is equal y on (3-(-2))= y on (3+2) =y on exponent5

jhonyy9 (jhonyy9):

and y5 squared is y10

OpenStudy (anonymous):

okay now I understand

OpenStudy (cathyangs):

\[(_{140x ^{3}y ^{-2}}^{28xy ^{3}})^{2}\] =\[(_{5x ^{2}y ^{-2}}^{1*1*y ^{3}}])^{2}\] =\[(_{5x ^{2}}^{y ^{5}})^{2}\] =\[_{25x ^{?}}^{y ^{10}}\]

jhonyy9 (jhonyy9):

all is sure ?

OpenStudy (anonymous):

yes I am sure and cathy thank you for the breakdown it is understandable.

jhonyy9 (jhonyy9):

so than good luck bye

OpenStudy (anonymous):

you too

OpenStudy (cathyangs):

you're talking to a newly-discovered girl who's now named Cathy...glad to help! :D

OpenStudy (anonymous):

XD Thank you will be posting more problems so be wary

OpenStudy (cathyangs):

ok! :D